40 grams of a 50% acid solution. How much water must be added to make a 10% solution
Answer by checkley77(12569)

.50*40=.10(40+x)
20=4+.10x
.10x=20-4
.10x=16
x=16/.10
x=160 grams of water is used.
Proof:
.50*40=.10(40+160)
20=.10*200
20=20